### v1

@@ VI Conclusion
We showed that an appropriate model uncertainty such as MUMI has the potential to reveal the mismatch between the training and test environments. However, since regression models equipped with MUMI are computationally complex, we resort to a classification paradigm, where we used a special label softening to tailor the method to localization. By using classification for localization, the model outputs reveal the peakwise uncertainty, which we then use to enhance the localization based on the source power.
@@
This paper shows that by suitable domain adaptation, one can enhance localization robustness in data-driven methods. With more prior knowledge about domain shifts between the training and testing environments and source location distribution, one may be able to improve the performance by using more effective methods for selecting $\mathcal{S}$ even without access to training data. However, the performance improvement will be limited in the absence of such knowledge. Future potential directions on this front include investigating other TTA methods currently used for other machine learning problems and developing other methods uncertainty quantification in the stated UWA problem.

### v2

@@ VII Conclusion
In this paper, we tackled test-time adaptation of a deep-learning based localization algorithm to environmental mismatch between training and test data. We considered different forms of mismatch including that in water depth, bathymetric shape and sediment type. We showed that an appropriate model uncertainty such as MUMI has the potential to reveal the mismatch between the training and test environments. Since regression models equipped with MUMI are computationally complex, we resort to a classification paradigm, using a special label softening to tailor the method to localization. By using classification for localization, the model outputs reveal the peakwise uncertainty, which we then use to enhance the localization based on the source power.
@@
This paper shows that by suitable domain adaptation, one can enhance localization robustness in data-driven methods. With more prior knowledge about domain shifts between the training and testing environments and source location distribution, one may be able to improve the performance by using more effective methods for selecting $\mathcal{S}$ even without access to training data. However, the performance improvement will be limited in the absence of such knowledge. Future potential directions on this front include extending the method to broadband sources or multiple sources, investigating other TTA methods currently used for other machine learning problems and developing other methods for uncertainty quantification in the stated UWA problem.
Appendix A Complexity Analysis
Figure 19: A typical convolutional layer with $H_{\text{in}}$ input channels and $H_{\text{out}}$ output channels.
Fig. 19 shows the size of tensors involved in a typical convolutional layer, where there are $H_{\text{in}}$ input channels, there are $H_{\text{out}}$ output channels, and the input and output dimensions are both $L\times L$ due to to padding. According to this figure, the number of multiplications in a convolutional layer is $H_{\text{in}}H_{\text{out}}L^{2}k^{2}$ and the number of additions is $H_{\text{in}}H_{\text{out}}L^{2}k^{2}$, resulting in a computational complexity of $2H_{\text{in}}H_{\text{out}}L^{2}k^{2}$. This yields a computational complexity of $O(L^{2})$ and a memory complexity of $H_{\text{in}}H_{\text{out}}L^{2}$. Moreover, according to Fig. 19, each convolutional layer require $H_{\text{in}}H_{\text{out}}k^{2}$ floating point numbers to be saved, while each linear layer requires a matrix of size input-size $\times$ output-size to be saved. Also, each layer requires an activation memory to hold the input and output tensors. Table IV summarizes the computational and memory complexities of each layer during a forward pass. In addition, updating the network weights using samples from the test environment involves backpropagation. Each backpropagation iteration requires the same amount of memory that a forward pass needs, to save the gradients with respect to inputs and weights. However, the computational complexity of a backpropagation pass is twice that of a forward pass, as backpropagation involves computing gradients with respect to not only the weights, but also inputs.
TABLE IV: CNN Architecture Complexity Analysis
Layer Type $H_{in}$ $H_{out}$ Input Size Output Size
Computational
Complexity
Memory
Complexity
1 conv (k=3) 2 6 $\times L L×L$$\times L$ $216L^{2}$ $2L^{2}+6L^{2}+108$
2 conv (k=5) 6 38 $\times L L×L$$\times L$ $11{,}400L^{2}$ $38L^{2}+5{,}700$
3 conv (k=5) 38 40 $\times L L×L$$\times L$ $76{,}000L^{2}$ $40L^{2}+38{,}000$
4 linear 11 11 $40L^{2}$ $N_{\phi}$ $80N_{\phi}L^{2}$ $40L^{2}N_{\phi}$
5 linear 11 11 $N_{\phi}$ $M$ $2N_{\phi}M$ $MN_{\phi}$
The JSEA methods refines the outputs of the CNN to achieve the final estimates for the source ranges. While the aforementioned complexity analysis applies to any general CNN architecture, the JSEA refinement step requires saving the estimates and energy levels of the received signals for a subset of the test samples (certain samples). Therefore, this refinement step increases the memory complexity by $O(N_{\text{test}})$ floating point numbers. Furthermore, peak finding algorithm may be performed with $O(MW)$ complexity, where $W$ is the size of the sliding window over which we find the local maximum and $M$ is the number of classes. Also, JSEA involves computing the received signal power, which adds to the complexity by $O(M)$ per sample. Finally, for an uncertain sample with $N_{P}$ significant peaks, we need to construct $\mathcal{S}_{\delta}(d)\subseteq\mathcal{S}$ for each peak and compare the average received power over that set to that of the uncertain sample, which can be done with a complexity of $O(N_{P}|\mathcal{S}|)$. Note that while JSEA refinement provides final range estimates, fine-tuning the model (model adaptation) using the pseudo-labeled certain samples is a necessary step in SHOT and involves backpropagation for $N_{\text{itr}}$ iterations. Since each backpropagation iteration is twice as complex as a forward CNN pass, this model adaptation step has a complexity of $2N_{\text{itr}}$ times that of the forward pass, as mentioned in Table V.
Matched field processing method in (3) involves two complex matrix multiplications. Since $\mathbf{\tilde{r}}(f,d)$ is of size $L\times 1$, $d\in\{d_{1},...,d_{N_{\text{tr}}}\}$, and $\tilde{\mathbf{C}}$ is of size $L\times L$, this yields a memory complexity of $2N_{\text{tr}}L$ floating point numbers and a computational complexity of $N_{\text{tr}}(8L^{2}+8L)$. Table V summarizes the complexities of different modules.
TABLE V: Memory and Computational Complexity Comparison
Model Computational Complexity Memory Complexity
MFP $N_{\text{tr}}(8L^{2}+8L)$ $2N_{\text{tr}}L$
CNN-c forward pass $(87,616+80N_{\phi})L^{2}+2MN_{\phi}$ $(86+40N_{\phi})L^{2}+MN_{\phi}+43,808$
JSEA-c refinement $O(M+MW+N_{P}|\mathcal{S}|)$ $N_{\text{test}}(L+1)$
JSEA-c or SHOT model adaptation $2N_{\text{itr}}((87,616+80N_{\phi})L^{2}+2N_{c}N_{\phi})$ $N_{\text{test}}L+(86+40N_{\phi})L^{2}+MN_{\phi}+43,808$